# Multi-Domain Evolutionary Optimization of Network Structures

### v1

@@ V-B Experiment on real-world networks
Figure 2: The evolution process of MDEO and SDEO. The attacker is FastGreedy.
Figure 3: The evolution process of MDEO and SDEO. The attacker is WalkTrap.
V-C Parameter sensitivity analysis
To bolster the completeness of our study, we conduct a targeted sensitivity analysis. In order to validate whether the transfer contributes to the optimization, we investigate the influence of the number of networks assisting in the target network on the results, as shown in Table II. For fairness, the number of solutions to transfer is the same for each setting. As seen, the results are related to the number of assisted networks, as reflected by that the fitness values of $|\mathcal{G}^{\mathcal{S}}|=2$ and $|\mathcal{G}^{\mathcal{S}}|=3$ are larger than that of $|\mathcal{G}^{\mathcal{S}}|=1$. Furthermore, in the test of a total of eight networks, $|\mathcal{G}^{\mathcal{S}}|=3$ achieves the optimum on five networks, five networks, and seven networks respectively when defending against FastGreedy and WalkTrap, meaning our transfer strategy is effective in the optimization.
We also examine the sensitivity of the ratio of elite solutions that will be directly transferred into the next generation in evolution. We refer this ratio as $\mu$ and set it to $\mu_{1}=0.05$, $\mu_{2}=0.10$, $\mu_{3}=0.15$ and $\mu_{4}=0.20$, respectively. We can observe that no ratio exhibits obvious advantages over other settings. For example, the large ratio of elite $\mu_{4}$ (purple line) achieves the best performance on Adjnoun, Dolphins, Lesmis, Erods and Polbooks while achieving the worst performance on USAir and BioCelegans. On the contrary, the setting of low ratio $\mu_{1}$, which has the best performance on USAir and BioCelegans, is inferior to other settings on the rest of networks. These two examples demonstrate that MDEO is not sensitive to the ratio of elite solutions.
TABLE II: The fitness comparison of the number of networks participating in transfer in the process of optimization.
Network FastGreedy WalkTrap
$|\mathcal{G}^{\mathcal{S}}|=1$ $|\mathcal{G}^{\mathcal{S}}|=2$ $|\mathcal{G}^{\mathcal{S}}|=3$ $|\mathcal{G}^{\mathcal{S}}|=1$ $|\mathcal{G}^{\mathcal{S}}|=2$ $|\mathcal{G}^{\mathcal{S}}|=3$
Dolphins 0.177 0.198 0.199 0.213 0.231 0.237
Lesmis 0.190 0.193 0.195 0.110 0.105 0.113
Polbooks 0.044 0.051 0.049 0.050 0.051 0.047
Adjnoun 0.684 0.653 0.649 0.302 0.297 0.312
Erdos 0.141 0.156 0.159 0.046 0.060 0.046
USAir 0.031 0.038 0.040 0.023 0.022 0.024
Netscience 0.031 0.038 0.039 0.041 0.040 0.052
BioCelegans 0.101 0.110 0.106 0.086 0.094 0.100
Figure 4: The fitness profile comparison of different elite transfer ratios.
V-D Structural change
@@
The structural change in some indices after perturbation is shown in Table III, through which we can observe which kind of edges that the community deception algorithm prefers to modify. CC refers to the clustering coefficient, measuring the extent to which nodes tend to cluster together. The clustering coefficients of all networks have decreased, meaning the local connectivity within the network is becoming lower. On the other hand, the average shortest distance (ASD) gets improved after rewiring edges, which suggests the distance between communities is shortened. As for the centralities, the change in Betweenness is universally larger than the change in PageRank, indicating the proposed community deception algorithm tends to change the bridge edges. The observed reduction in modularity indicates that our algorithm effectively weakens the internal connections of individual communities while simultaneously enhancing their external connectivity.
TABLE III: The structural change after modifying edges obtained by MDEO-II. The tested attacker is FastGreedy.
Data $|E^{+}\cup E^{-}|$ CC ASD $20\%-$ Betweenness $20\%-$ PageRank Modularity
@@
VI Conclusion
In this work, we have explored a framework–multi-domain evolutionary optimization. We actualized this concept within the realm of network structures, predicated on the inherent shared properties of real-world networks. The proposed MDEO consists of three components: 1) measuring the similarity of domains at the community level; 2) selecting the anchor pair nodes and training the mapping of solutions of different domains 3) optimizing different domains simultaneously and making them promote each other. The proposed method was validated on the task of community deception with eight different real-world networks of varying sizes from various domains, and the experimental results (with statistical analysis and sensitivity analysis) show the fitness values of evolutionary optimization on different networks have been improved, suggesting the effectiveness of our proposed multi-domain evolutionary optimization. As an exploratory work, MDEO has been successfully applied to networks of similar sizes. The knowledge transfer from a large network to a small network or vice versa, has not been studied yet. In future work, we will develop a more robust and adaptive framework that allows networks of various scales to be optimized simultaneously with improved effectiveness.

### v2

@@ V-C Experiment on real-world networks
Figure 2: The optimization curve of MDEO and SDEO on eight tested networks for the illustrative task of community deception, with FastGreedy as the attacker.
Figure 3: The optimization curve of MDEO and SDEO on eight tested networks for the illustrative task of community deception, with WalkTrap as the attacker.
V-D Parameter sensitivity analysis
In this section, we investigate the parameter sensitivity of the MDEO by varying three key parameters in Table IV: The transfer interval (k), the number of transfer solutions ($|\text{T}|$), and the number of assisted networks ($|\mathcal{G}^{S}|$), with the default configuration set to $\text{k}=5$, $|\text{T}|=30$, and $|\mathcal{G}^{S}|=3$. Compared to the default setting, increasing the transfer interval (e.g., $\text{k}=10$ and $\text{k}=20$) generally results in slightly degraded performance across most networks, indicating that more frequent knowledge transfer (lower k) is beneficial for optimization. Likewise, reducing the number of transfer solutions (e.g., $|\text{T}|=10$ and $|\text{T}|=20$) leads to small drops in performance when compared with the default $|\text{T}|=30$, suggesting that transferring more solutions helps guide the search more effectively. Furthermore, reducing the number of assisted networks (i.e., $|\mathcal{G}^{S}|=1$ or $2$) also yields worse results relative to the default, reinforcing the idea that leveraging more source networks provides richer and more useful guidance. We also investigate the role of community partitioning in network alignment by comparing community-based alignment with direct alignment of high- and low-degree nodes. The results show that community-based alignment more effectively facilitates knowledge transfer. The results clearly support that more frequent and comprehensive knowledge transfer leads to better optimization outcomes, indicating the effectiveness of the proposed MDEO.
TABLE IV: Parameter sensitivity analysis of MDEO with the knowledge transfer interval k set to 5, the number of transfer solutions $|\text{T}|$ set to 30, and the number of assisted networks $|\mathcal{G}^{\mathcal{S}}|$ set to 3. MDEO incorporates community-based alignment by default, referred to CA. The tested attacker is the FastGreedy algorithm. The shadow cell is a better setting than default.
  Network W/O CA $\mathbf{k=10}$ $\mathbf{k=20}$ $\mathbf{|T|=10}$ $\mathbf{|T|=20}$ $\boldsymbol{|\mathcal{G}^{\mathcal{S}}|=1}$ $\boldsymbol{|\mathcal{G}^{\mathcal{S}}|=2}$ Default
Dolphins 0.189 0.187 0.182 0.188 0.189 0.167 0.188 0.198
Lesmis 0.213 0.195 0.186 0.182 0.197 0.205 0.207 0.208
Polbooks 0.044 0.048 0.041 0.038 0.047 0.044 0.051 0.051
Adjnoun 0.644 0.623 0.621 0.578 0.633 0.644 0.653 0.649
Erdos 0.151 0.149 0.150 0.146 0.149 0.141 0.148 0.152
USAir 0.029 0.032 0.028 0.028 0.037 0.030 0.037 0.039
Netscience 0.028 0.030 0.027 0.030 0.032 0.029 0.031 0.032
BioCelegans 0.104 0.106 0.104 0.105 0.106 0.101 0.105 0.108
Comparison 1/7 0/8 0/8 0/8 0/8 0/8 1/7 -
As for the efficiency, Figure 4 illustrates the average running time of the MDEO for different parameter configurations related to knowledge transfer. The left subplot examines the impact of varying the transfer interval (k) while the right subplot evaluates different transfer solution numbers ($|\text{T}|$). The results indicate that both parameters have a marginal influence on the average running time. Specifically, as the transfer interval increases, a slight decrease in running time is observed, suggesting reduced overhead from less frequent transfers. Conversely, increasing the number of transfer solutions slightly increases the running time, likely due to the additional computational cost of processing more solutions. As observed, the impact of knowledge transfer settings on running time is minimal, indicating that the MDEO achieves an excellent balance between effectiveness and efficiency.
Figure 4: Average running time of MDEO under different parameter settings when optimizing eight real-world networks simultaneously.
V-E Structural change
@@
The structural change in some indices after perturbation is shown in Table V, through which we can observe which kind of edges that the community deception algorithm prefers to modify. CC refers to the clustering coefficient, measuring the extent to which nodes tend to cluster together. The clustering coefficients of all networks have decreased, meaning the local connectivity within the network is becoming lower. On the other hand, the average shortest distance (ASD) gets improved after rewiring edges, which suggests the distance between communities is shortened. As for the centralities, the change in Betweenness is universally larger than the change in PageRank, indicating the proposed community deception algorithm tends to change the bridge edges. The observed reduction in modularity indicates that our algorithm effectively weakens the internal connections of individual communities while simultaneously enhancing their external connectivity.
TABLE V: The structural modifications resulting from edge alterations performed by MDEO, with FastGreedy as the testing attacker.
  Data $|E^{+}\cup E^{-}|$ $C$ ASD $20\%$- Betweenness $20\%$- PageRank Modularity
@@
V-F Generalizability analysis
The MDEO framework, originally demonstrated through the edge-level task of community deception, is fundamentally designed for generalizability. It can be effectively applied to a wide range of graph-structured combinatorial problems, encompassing both node-level and edge-level objectives. By appropriately adapting the solution representation and fitness function to suit the specific optimization goal, MDEO is capable of addressing diverse challenges across various domains.
As a way of illustration, we test the classical node-level task, i.e., the influence maximization task [67], where the solution sizes are set to 10 for the Email and Polblogs networks, and 15 for the Facebook and Wiki networks. As shown in Figures 5, MDEO consistently achieves higher average fitness values compared to SDEO, highlighting its effectiveness across different network structures. These representative tasks at each level collectively demonstrate the method’s generality and adaptability.
Figure 5: The optimization curve of MDEO and SDEO in addressing the node-level problem of influence maximization.
VI Discussion
Although our MDEO framework has demonstrated promising performance, there still exists some challenges. Many real-world networks vary considerably in scale, and direct knowledge transfer across such imbalanced scenarios may introduce negative transfer effects. To address this issue, we observe that certain fundamental structural properties, such as modularity, centrality, and connectivity patterns, tend to remain consistent across different scales, providing a reliable foundation for effective knowledge transfer between networks of varying sizes. Building on this insight, we will explore extending MDEO to support a bidirectional transfer mechanism that enables adaptive knowledge exchange between large and small networks. This enhancement is designed to improve the robustness and adaptability of our approach in imbalanced-size scenarios, thereby enhancing its effectiveness across heterogeneous domains. Developing this bidirectional optimization framework is a key direction for future work and is expected to expand the practical applicability of MDEO.
VII Conclusion
In this work, we have explored a framework–multi-domain evolutionary optimization (MDEO). We actualized this concept within the realm of network structures, predicated on the inherent shared properties of real-world networks. The proposed method was validated on the task of community deception with eight different real-world networks of varying sizes from various domains, and the experimental results show that the fitness values of evolutionary optimization on different networks have been improved, suggesting the effectiveness of our proposed MDEO. As an exploratory work, MDEO has been successfully applied to networks of similar sizes. The knowledge transfer from a large network to a small network or vice versa, has not been studied yet. In future work, we will develop a more robust and adaptive framework that allows networks of various scales to be optimized simultaneously with improved effectiveness.